A group of 5 patients treated with Medicine type A weight 42, 39, 48, 60 and 41 kg.


A second group of 5 patients treated with Medicine type B weight 38, 42, 48, 67, 40 kg. Do


the two medicines differ significantly with regard to their effect and increasing weight? [At


5% level of significance]

1
Expert's answer
2022-01-27T12:25:30-0500
"\bar{x}_1=\dfrac{42+39+48+60+41}{5}=46""s_1^2=\dfrac{1}{5-1}((42-46)^2+(39-46)^2+(48-46)^2""+(60-46)^2+(41-46)^2)=\dfrac{145}{2}"




"s_1=\sqrt{s_1^2}=\sqrt{\dfrac{145}{2}}\approx8.5147"





"\bar{x}_2=\dfrac{38+42+48+67+40}{5}=47""s_2^2=\dfrac{1}{5-1}((38-47)^2+(42-47)^2+(48-47)^2""+(67-47)^2+(40-47)^2)=139"




"s_2=\sqrt{s_2^2}=\sqrt{139}\approx11.7898"


A F-test is used to test for the equality of variances. The following F-ratio is obtained:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{72.5}{139}=0.5216"

The critical values for two-tailed, "df_1=df_2=5-1=4" are "F_L = 0.1041"  and "F_U = 9.6045," and since "F = 0.5216," then the null hypothesis of equal variances is not rejected.

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

The significance level is "\alpha = 0.05," and the degrees of freedom are "df =n_1+n_2-2= 5+5-2=8"

Hence, it is found that the critical value for this two-tailed test, "\alpha = 0.05" and "df = 8" is "t_c = 2.3060."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.3060\}."

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t=\dfrac{\bar{x}_1-\bar{x}_2}{\sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}(\dfrac{1}{n_1}+\dfrac{1}{n_2})}}"

"=\dfrac{46-47}{\sqrt{\dfrac{(5-1)(72.5)+(5-1)(139)}{5+5-2}(\dfrac{1}{5}+\dfrac{1}{5})}}"

"=-0.153755"

Since it is observed that "|t| = 0.153755 \le 2.3060=t_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=8" degrees of freedom, "t=-0.153755," is "p=0.881611," and since "p = 0.881611 \ge 0.05=\alpha," it is concluded that the null hypothes is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha = 0.05" significance level.